# LoCoDL: Communication-Efficient Distributed Learning with Local Training and Compression

### v1

@@ 1 Introduction
Performing distributed computations is now pervasive in all areas of science. Notably, Federated Learning (FL) consists in training machine learning models in a distributed and collaborative way (Konečný et al., 2016a, b; McMahan et al., 2017; Bonawitz et al., 2017). The key idea in this rapidly growing field is to exploit the wealth of information stored on distant devices, such as mobile phones or hospital workstations. The many challenges to face in FL include data privacy and robustness to adversarial attacks, but communication-efficiency is likely to be the most critical (Kairouz et al., 2021; Li et al., 2020a; Wang et al., 2021). Indeed, in contrast to the centralized setting in a datacenter, in FL the clients perform parallel computations but also communicate back and forth with a distant orchestrating server. Communication typically takes place over the internet or cell phone network, and can be slow, costly, and unreliable. It is the main bottleneck that currently prevents large-scale deployment of FL in mass-market applications.
@@ 1.1 Problem and Motivation
where $d\geq 1$ is the model dimension and the functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are smooth, so their gradients will be called. We consider the server-client model in which $n\geq 1$ clients do computations in parallel and communicate back and forth with a server. The private function $f_{i}$ is owned by and stored on client $i\in[n]\coloneqq\{1,\ldots,n\}$. Problem (1) models empirical risk minimization, of utmost importance in machine learning (Sra et al., 2011; Shalev-Shwartz & Ben-David, 2014). More generally, minimizing a sum of functions appears in virtually all areas of science and engineering. Our goal is to solve Problem (1) in a communication-efficient way, in the general heterogeneous setting in which the functions $f_{i}$, as well as $g$, can be arbitrarily different: we do not make any assumption on their similarity whatsoever.
We consider in this work the strongly convex setting — an analysis with nonconvex functions would certainly require very different proof techniques, which we currently do not know how to derive. That is, the following holds:
@@
  • Viewpoint 1: regularization. The function $g$ can be a regularizer. For instance, if the functions $f_{i}$ are convex, adding $g=\frac{\mu}{2}\|\cdot\|^{2}$ for a small $\mu>0$ makes the problem $\mu$-strongly convex.
  • Viewpoint 2: shared dataset. The function $g$ can model the cost of a common dataset, or a piece thereof, that is known to all clients.
  • Viewpoint 3: server-aided training. The function $g$ can model the cost of a core dataset, known only to the server, which makes calls to $\nabla g$. This setting has been investigated in several works, with the idea that using a small auxiliary dataset representative of the global data distribution, the server can correct for the deviation induced by partial participation (Zhao et al., 2018; Yang et al., 2021, 2023). We do not focus on this setting, because we deal with the general heterogeneous setting in which $g$ and the $f_{i}$ are not meant to be similar in any sense, and in our work $g$ is handled by the clients, not by the server.
  • Viewpoint 4: a new mathematical and algorithmic principle. This is the idea that led to the construction of LoCoDL, and we detail it in Section 2.1.
@@ 1.2 State of the Art
We review the latest developments on communication-efficient algorithms for distributed learning, making use of LT, CC, or both. Before that, we note that we should distinguish uplink, or clients-to-server, from downlink, or server-to-clients, communication. Uplink is usually slower than downlink communication, since uploading different messages in parallel to the server is slower than broadcasting the same message to an arbitrary number of clients. This can be due to cache memory and aggregation speed constraints of the server, as well as asymmetry of the service provider’s systems or protocols used on the internet or cell phone network. In this work, we focus on the uplink communication complexity, which is the bottleneck in practice. Indeed, the goal is to exploit parallelism to obtain better performance when $n$ increases. Precisely, with LoCoDL, the uplink communication complexity decreases from $\mathcal{O}\left(d\sqrt{\kappa}\log\epsilon^{-1}\right)$ when $n$ is small to $\mathcal{O}\big{(}\sqrt{d}\sqrt{\kappa}\log\epsilon^{-1}\big{)}$ when $n$ is large, where the condition number $\kappa$ is defined in Assumption 1.1, see Corollary 3.2. Many works have considered bidirectional compression, which consists in compressing the messages sent both ways (Gorbunov et al., 2020b; Philippenko & Dieuleveut, 2020; Liu et al., 2020; Philippenko & Dieuleveut, 2021; Condat & Richtárik, 2022; Gruntkowska et al., 2023; Tyurin & Richtárik, 2023b) but to the best of our knowledge, this has no impact on the downlink complexity, which cannot be reduced further than $\mathcal{O}\left(d\sqrt{\kappa}\log\epsilon^{-1}\right)$, just because there is no parallelism to exploit in this direction. Thus, we focus our analysis on theoretical and algorithmic techniques to reduce the uplink communication complexity, which we call communication complexity in short, and we ignore downlink communication.
@@ 1.3 A General Class of Unbiased Random Compressors
For every $\omega\geq 0$, we define the $\mathbb{U}(\omega)$ as the set of random compression operators $\mathcal{C}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that are unbiased, i.e. $\mathbb{E}\!\left[\mathcal{C}(x)\right]=x$, and satisfy, for every $x\in\mathbb{R}^{d}$,
@@ 1.4 Challenge and Contributions
This work addresses the following question:
Can we combine LT and CC with any compressors in the generic class $\mathbb{U}(\omega)$ defined in the previous section, and fully benefit from both techniques by obtaining a doubly-accelerated communication complexity?
We answer this question in the affirmative. LoCoDL has the same probabilistic LT mechanism as Scaffnew and features CC with compressors in $\mathbb{U}(\omega)$ with arbitrarily large $\omega\geq 0$, with proved linear convergence under Assumption 1.1, without further requirements. By choosing the communication probability and the variance $\omega$ appropriately, double acceleration is obtained. Thus, LoCoDL achieves the same theoretical complexity as CompressedScaffnew, but allows for a large class of compressors instead of the cumbersome permutation-based compressor of the latter. In particular, with compressors performing sparsification and quantization, LoCoDL outperforms existing algorithms, as we show by experiments in Section 4. This is remarkable, since ADIANA, based on Nesterov acceleration and not LT, has an even better theoretical complexity when $n$ is larger than $d$, see Table 2, but this is not reflected in practice: ADIANA is clearly behind LoCoDL in our experiments. Thus, LoCoDL sets new standards in terms of communication efficiency.
@@
[…]
@@ 2.1 Principle: Double Lifting of the Problem to a Consensus Problem
Local steps with local models.
In algorithms making use of LT, such as FedAvg, Scaffold and Scaffnew, the clients store and update local model estimates $x_{i}$. When communication occurs, an estimate of their average is formed by the server and broadcast to all clients. They all resume their computations with this new model estimate.
Compressing the difference between two estimates.
To implement CC, a powerful idea is to compress not the vectors themselves, but difference vectors that converge to zero. This way, the algorithm is variance-reduced; that is, the compression error vanishes at convergence. The technique of compressing the difference between a gradient vector and a control variate is at the core of algorithms such as DIANA and EF21. Here, we want to compress differences between model estimates, not gradient estimates. That is, we want Client $i$ to compress the difference between $x_{i}$ and another model estimate that converges to the solution $x^{\star}$ as well. We see the need of an additional model estimate that plays the role of an anchor for compression. This is the variable $y$ common to all clients in LoCoDL, which compress $x_{i}-y$ and send these compressed differences to the server.
Combining the two ideas.
Accordingly, an equivalent reformulation of (1) is the consensus problem with $n+1$ variables
@@
The primal–dual optimality conditions are
$x_{1}=\cdots=x_{n}=y,$
$0=\nabla f_{i}(x_{i})-u_{i},\ \forall i\in[n],$
$0=\nabla g(y)-v,$
$0=u_{1}+\cdots+u_{n}+nv,$ (4)
for some dual variables $u_{1},\ldots,u_{n},v$ introduced in LoCoDL, that always satisfy the condition (4) of dual feasibility.
@@
[…]
@@
16:           $x_{i}^{t+1}\coloneqq\hat{x}_{i}^{t}$, $y^{t+1}=\hat{y}^{t}$
17:           $u_{i}^{t+1}\coloneqq u_{i}^{t}$, $v^{t+1}\coloneqq v^{t}$
18:        end if
@@ 2.2 Description of LoCoDL
LoCoDL is a randomized primal–dual algorithm, shown as Algorithm 1. At every iteration, for every $i\in[n]$ in parallel, Client $i$ first constructs a prediction $\hat{x}_{i}^{t}$ of its updated local model estimate, using a GD step with respect to $f_{i}$ corrected by the dual variable $u_{i}^{t}$. It also constructs a prediction $\hat{y}^{t}$ of the updated model estimate, using a GD step with respect to $g$ corrected by the dual variable $v^{t}$. Since $g$ is known by all clients, they all maintain and update identical copies of the variables $y$ and $v$. If there is no communication, which is the case with probability $1-p$, $x_{i}$ and $y$ are updated with these predicted estimates, and the dual variables $u_{i}$ and $v$ are unchanged. If communication occurs, which is the case with probability $p$, the clients compress the differences $\hat{x}_{i}^{t}-\hat{y}^{t}$ and send these compressed vectors to the server, which forms $\bar{d}^{t}$ equal to one half of their average. Then the variables $x_{i}$ are updated using a convex combination of the local predicted estimates $\hat{x}_{i}^{t}$ and the global but noisy estimate $\hat{y}^{t}+\bar{d}^{t}$. $y$ is updated similarly. Finally, the dual variables are updated using the compressed differences minus their weighted average, so that the dual feasibility (4) remains satisfied. The model estimates $x_{i}^{t}$, $\hat{x}_{i}^{t}$, $y^{t}$, $\hat{y}^{t}$ all converge to $x^{\star}$, so that their differences, as well as the compressed differences as a consequence of (2), converge to zero. This is the key property that makes the algorithm variance-reduced. We consider the following assumption.
@@ 3 Convergence and Complexity of LoCoDL
$\Psi^{t}$ $\coloneqq\frac{1}{\gamma}\left(\sum_{i=1}^{n}\left\|x_{i}^{t}-x^{\star}\right\|^{2}+n\left\|y^{t}-x^{\star}\right\|^{2}\right)$ (5)
$\quad+\frac{\gamma(1+2\omega)}{p^{2}\chi}\left(\sum_{i=1}^{n}\left\|u_{i}^{t}-u_{i}^{\star}\right\|^{2}+n\left\|v^{t}-v^{\star}\right\|^{2}\right),$
@@
$\mathbb{E}\!\left[\Psi^{t}\right]\leq\tau^{t}\Psi^{0},$ (6)
where
$\tau\coloneqq\max\left((1-\gamma\mu)^{2},(1-\gamma L)^{2},1-\frac{p^{2}\chi}{1+2\omega}\right)<1.$ (7)
@@
The proof is derived in Appendix A.
We place ourselves in the conditions of Theorem 3.1. We observe that in (7), the larger $\chi$, the better, so given $\rho$ we should set
$\chi=2\rho-\rho^{2}(1+\omega_{\mathrm{av}}).$ (8)
Then, choosing $\rho$ to maximize $\chi$ yields
@@
We now study the complexity of LoCoDL with $\chi$ and $\rho$ chosen as in (9) and $\gamma=\Theta(\frac{1}{L})$.
We can remark that LoCoDL has the same rate $\tau^{\sharp}\coloneqq\max(1-\gamma\mu,\gamma L-1)^{2}$ as mere distributed gradient descent, as long as $p^{-1}$, $\omega$ and $\omega_{\mathrm{av}}$ are small enough to have $1-\frac{p^{2}\chi}{1+2\omega}\leq\tau^{\sharp}$. This is remarkable: communicating with a low frequency and compressed vectors does not harm convergence at all, until some threshold.
@@
the iteration complexity becomes
$\mathcal{O}\Big{(}\big{(}\kappa+\omega(1+\omega_{\mathrm{av}})\big{)}\log\epsilon^{-1}\Big{)}$
and the communication complexity in number of communication rounds is $p$ times the iteration complexity, that is
@@
Let us consider the example of independent rand-$k$ compressors, for some $k\in[d]$. We have $\omega=\frac{d}{k}-1$. Therefore, the communication complexity in numbers of reals is $k$ times the complexity in number of rounds; that is,
$\mathcal{O}\left(\left(\left(\sqrt{kd}+\frac{d}{\sqrt{n}}\right)\sqrt{\kappa}+d\left(1+\frac{d}{kn}\right)\right)\log\epsilon^{-1}\right).$
We can now choose $k$ to minimize this complexity: with $k=\lceil\frac{d}{n}\rceil$, it becomes
$\mathcal{O}\left(\left(\left(\sqrt{d}+\frac{d}{\sqrt{n}}\right)\sqrt{\kappa}+d\right)\log\epsilon^{-1}\right),$
as shown in Table 2. Let us state this result:
@@
  1. 1.
    In the “large $d$ small $n$” regime, i.e. $n=\mathcal{O}(d)$, the communication complexity of LoCoDL in (13) becomes
    $\mathcal{O}\left(\left(\frac{d\sqrt{\kappa}}{\sqrt{n}}+d\right)\log\epsilon^{-1}\right).$
    This is the state of the art, as reported in Table 2.
  2. 2.
    In the “large $n$ small $d$” regime, i.e. $n=\Omega(d)$, the communication complexity of LoCoDL in (13) becomes
    $\mathcal{O}\left(\left(\sqrt{d}\sqrt{\kappa}+d\right)\log\epsilon^{-1}\right).$
    If $n$ is even larger with $n=\Omega(d^{2})$, ADIANA achieves the even better complexity
    $\mathcal{O}\left(\left(\sqrt{\kappa}+d\right)\log\epsilon^{-1}\right).$

### v2

@@ 1 Introduction
Performing distributed computations is now pervasive in all areas of science. Notably, Federated Learning (FL) consists in training machine learning models in a distributed and collaborative way (Konečný et al., 2016a; b; McMahan et al., 2017; Bonawitz et al., 2017). The key idea in this rapidly growing field is to exploit the wealth of information stored on distant devices, such as mobile phones or hospital workstations. The many challenges to face in FL include data privacy and robustness to adversarial attacks, but communication-efficiency is likely to be the most critical (Kairouz et al., 2021; Li et al., 2020a; Wang et al., 2021). Indeed, in contrast to the centralized setting in a datacenter, in FL the clients perform parallel computations but also communicate back and forth with a distant orchestrating server. Communication typically takes place over the internet or cell phone network, and can be slow, costly, and unreliable. It is the main bottleneck that currently prevents large-scale deployment of FL in mass-market applications.
@@ 1.1 Problem and Motivation
where $d\geq 1$ is the model dimension and the functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are smooth. We consider the server-client model in which $n\geq 1$ clients do computations in parallel and communicate back and forth with a server. The private function $f_{i}$ is owned by and stored on client $i\in[n]\coloneqq\{1,\ldots,n\}$. Problem (1) models empirical risk minimization, of utmost importance in machine learning (Sra et al., 2011; Shalev-Shwartz & Ben-David, 2014). More generally, minimizing a sum of functions appears in virtually all areas of science and engineering. Our goal is to solve Problem (1) in a communication-efficient way, in the general heterogeneous setting in which the functions $f_{i}$, as well as $g$, can be arbitrarily different: we do not make any assumption on their similarity whatsoever.
We consider in this work the strongly convex setting. That is, the following holds:
@@
[…]
@@ 1.2 State of the Art
We review the latest developments on communication-efficient algorithms for distributed learning, making use of LT, CC, or both. Before that, we note that we should distinguish uplink, or clients-to-server, from downlink, or server-to-clients, communication. Uplink is usually slower than downlink communication, since the clients uploading different messages in parallel to the server is slower than the clients downloading the same message in parallel from the server. This can be due to cache memory and aggregation speed constraints of the server, as well as asymmetry of the service provider’s systems or protocols used on the internet or cell phone network. In this work, we focus on the uplink communication complexity, which is often the bottleneck in practice. Indeed, the goal is to exploit parallelism to obtain better performance when $n$ increases. Precisely, with LoCoDL, the uplink communication complexity decreases from $\mathcal{O}\left(d\sqrt{\kappa}\log\epsilon^{-1}\right)$ when $n$ is small to $\mathcal{O}\big{(}\sqrt{d}\sqrt{\kappa}\log\epsilon^{-1}\big{)}$ when $n$ is large, where the condition number $\kappa$ is defined in Assumption 1.1, see Corollary 3.2. Many works have considered bidirectional compression, which consists in compressing the messages sent both ways (Gorbunov et al., 2020b; Philippenko & Dieuleveut, 2020; Liu et al., 2020; Philippenko & Dieuleveut, 2021; Condat & Richtárik, 2022; Gruntkowska et al., 2023; Tyurin & Richtárik, 2023b) but to the best of our knowledge, this has no impact on the downlink complexity, which cannot be reduced further than $\mathcal{O}\left(d\sqrt{\kappa}\log\epsilon^{-1}\right)$, just because there is no parallelism to exploit in this direction. Thus, we focus our analysis on theoretical and algorithmic techniques to reduce the uplink communication complexity, which we call communication complexity in short, and we ignore downlink communication.
@@ 1.3 A General Class of Unbiased Random Compressors
For every $\omega\geq 0$, we define $\mathbb{U}(\omega)$ as the set of random compression operators $\mathcal{C}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that are unbiased, i.e. $\mathbb{E}\!\left[\mathcal{C}(x)\right]=x$, and satisfy, for every $x\in\mathbb{R}^{d}$,
@@ 1.4 Challenge and Contributions
This work addresses the following question: Can we combine LT and CC with any compressors in the generic class $\mathbb{U}(\omega)$ defined in the previous section, and fully benefit from both techniques by obtaining a doubly-accelerated communication complexity?
We answer this question in the affirmative. LoCoDL has the same probabilistic LT mechanism as Scaffnew and features CC with compressors in $\mathbb{U}(\omega)$ with arbitrarily large $\omega\geq 0$, with proved linear convergence under Assumption 1.1, without further requirements. By choosing the communication probability and the variance $\omega$ appropriately, double acceleration is obtained. Thus, LoCoDL achieves the same theoretical complexity as CompressedScaffnew, but allows for a large class of compressors instead of the cumbersome permutation-based compressor of the latter. In particular, with compressors performing sparsification and quantization, LoCoDL outperforms existing algorithms, as we show by experiments in Section 4. This is remarkable, since ADIANA, based on Nesterov acceleration and not LT, has an even better theoretical complexity when $n$ is larger than $d$, see Table 2, but this is not reflected in practice: ADIANA is clearly behind LoCoDL in our experiments. Thus, our experiments indicate that LoCoDL sets new standards in terms of communication efficiency.
@@
lower bound2 $\left(1+\frac{\omega}{\sqrt{n}}\right)\!\sqrt{\kappa}+\omega$ $\left(1+\frac{\omega}{\sqrt{n}}\right)\!\sqrt{\kappa}+\omega$ $\left(1\!+\!\sqrt{\omega}\right)\!\sqrt{\kappa}+\omega$
@@ 2.1 Principle: Double Lifting of the Problem to a Consensus Problem
Local steps with local models. In algorithms making use of LT, such as FedAvg, Scaffold and Scaffnew, the clients store and update local model estimates $x_{i}$. When communication occurs, an estimate of their average is formed by the server and broadcast to all clients. They all resume their computations with this new model estimate.
Compressing the difference between two estimates.  To implement CC, a powerful idea is to compress not the vectors themselves, but difference vectors that converge to zero. This way, the algorithm is variance-reduced; that is, the compression error vanishes at convergence. The technique of compressing the difference between a gradient vector and a control variate is at the core of algorithms such as DIANA and EF21. Here, we want to compress differences between model estimates, not gradient estimates. That is, we want Client $i$ to compress the difference between $x_{i}$ and another model estimate that converges to the solution $x^{\star}$ as well. We see the need of an additional model estimate that plays the role of an anchor for compression. This is the variable $y$ common to all clients in LoCoDL, which compress $x_{i}-y$ and send these compressed differences to the server.
Combining the two ideas.  Accordingly, an equivalent reformulation of (1) is the consensus problem with $n+1$ variables
@@
The primal–dual optimality conditions are $x_{1}=\cdots=x_{n}=y$, $0=\nabla f_{i}(x_{i})-u_{i}\ \forall i\in[n]$, $0=\nabla g(y)-v$, and $0=u_{1}+\cdots+u_{n}+nv$ (dual feasibility), for some dual variables $u_{1},\ldots,u_{n},v$ introduced in LoCoDL, that always satisfy the dual feasibility condition.
@@
10:           at server: aggregate $\bar{d}^{t}\coloneqq\frac{1}{2n}\sum_{j=1}^{n}d_{j}^{t}$ and broadcast $\bar{d}^{t}$ to all clients
@@
16:           $x_{i}^{t+1}\coloneqq\hat{x}_{i}^{t}$, $y^{t+1}=\hat{y}^{t}$, $u_{i}^{t+1}\coloneqq u_{i}^{t}$, $v^{t+1}\coloneqq v^{t}$
17:        end if
18:     end for
@@ 2.2 Description of LoCoDL
LoCoDL is a randomized primal–dual algorithm, shown as Algorithm 1. At every iteration, for every $i\in[n]$ in parallel, Client $i$ first constructs a prediction $\hat{x}_{i}^{t}$ of its updated local model estimate, using a GD step with respect to $f_{i}$ corrected by the dual variable $u_{i}^{t}$. It also constructs a prediction $\hat{y}^{t}$ of the updated model estimate, using a GD step with respect to $g$ corrected by the dual variable $v^{t}$. Since $g$ is known by all clients, they all maintain and update identical copies of the variables $y$ and $v$. If there is no communication, which is the case with probability $1-p$, $x_{i}$ and $y$ are updated with these predicted estimates, and the dual variables $u_{i}$ and $v$ are unchanged. If communication occurs, which is the case with probability $p$, the clients compress the differences $\hat{x}_{i}^{t}-\hat{y}^{t}$ and send these compressed vectors to the server, which forms $\bar{d}^{t}$ equal to one half of their average. Then the variables $x_{i}$ are updated using a convex combination of the local predicted estimates $\hat{x}_{i}^{t}$ and the global but noisy estimate $\hat{y}^{t}+\bar{d}^{t}$. $y$ is updated similarly. Finally, the dual variables are updated using the compressed differences minus their weighted average, so that the dual feasibility condition remains satisfied. The model estimates $x_{i}^{t}$, $\hat{x}_{i}^{t}$, $y^{t}$, $\hat{y}^{t}$ all converge to $x^{\star}$, so that their differences, as well as the compressed differences as a consequence of (2), converge to zero. This is the key property that makes the algorithm variance-reduced. We consider the following assumption.
@@ 3 Convergence and Complexity of LoCoDL
$\Psi^{t}\coloneqq\frac{1}{\gamma}\left(\sum_{i=1}^{n}\left\|x_{i}^{t}-x^{\star}\right\|^{2}+n\left\|y^{t}-x^{\star}\right\|^{2}\right)+\frac{\gamma(1+2\omega)}{p^{2}\chi}\left(\sum_{i=1}^{n}\left\|u_{i}^{t}-u_{i}^{\star}\right\|^{2}+n\left\|v^{t}-v^{\star}\right\|^{2}\right),$ (4)
@@
$\mathbb{E}\!\left[\Psi^{t}\right]\leq\tau^{t}\Psi^{0},\quad\mbox{where}\quad\tau\coloneqq\max\left((1-\gamma\mu)^{2},(1-\gamma L)^{2},1-\frac{p^{2}\chi}{1+2\omega}\right)<1.$ (5)
@@
We place ourselves in the conditions of Theorem 3.1. We observe that in (5), the larger $\chi$, the better, so given $\rho$ we should set $\chi=2\rho-\rho^{2}(1+\omega_{\mathrm{av}})$. Then, choosing $\rho$ to maximize $\chi$ yields
@@
We now study the complexity of LoCoDL with $\chi$ and $\rho$ chosen as in (6) and $\gamma=\Theta(\frac{1}{L})$. We remark that LoCoDL has the same rate $\tau^{\sharp}\coloneqq\max(1-\gamma\mu,\gamma L-1)^{2}$ as mere distributed gradient descent, as long as $p^{-1}$, $\omega$ and $\omega_{\mathrm{av}}$ are small enough to have $1-\frac{p^{2}\chi}{1+2\omega}\leq\tau^{\sharp}$. This is remarkable: communicating with a low frequency and compressed vectors does not harm convergence at all, until some threshold.
@@
the iteration complexity becomes $\mathcal{O}\Big{(}\big{(}\kappa+\omega(1+\omega_{\mathrm{av}})\big{)}\log\epsilon^{-1}\Big{)}$ and the communication complexity in number of communication rounds is $p$ times the iteration complexity, that is
@@
Let us consider the example of independent rand-$k$ compressors, for some $k\in[d]$. We have $\omega=\frac{d}{k}-1$. Therefore, the communication complexity in numbers of reals is $k$ times the complexity in number of rounds; that is, $\mathcal{O}\left(\left(\left(\sqrt{kd}+\frac{d}{\sqrt{n}}\right)\sqrt{\kappa}+d\left(1+\frac{d}{kn}\right)\right)\log\epsilon^{-1}\right)$. We can now choose $k$ to minimize this complexity: with $k=\lceil\frac{d}{n}\rceil$, it becomes $\mathcal{O}\left(\left(\left(\sqrt{d}+\frac{d}{\sqrt{n}}\right)\sqrt{\kappa}+d\right)\log\epsilon^{-1}\right)$, as shown in Table 2. Let us state this result:
@@
1. In the “large $d$ small $n$” regime, i.e. $n=\mathcal{O}(d)$, the communication complexity of LoCoDL in (10) becomes $\mathcal{O}\left(\left(\frac{d\sqrt{\kappa}}{\sqrt{n}}+d\right)\log\epsilon^{-1}\right)$. This is the state of the art, as reported in Table 2.
2. In the “large $n$ small $d$” regime, i.e. $n=\Omega(d)$, the communication complexity of LoCoDL in (10) becomes $\mathcal{O}\left(\left(\sqrt{d}\sqrt{\kappa}+d\right)\log\epsilon^{-1}\right)$. If $n$ is even larger with $n=\Omega(d^{2})$, ADIANA achieves the even better complexity $\mathcal{O}\left(\left(\sqrt{\kappa}+d\right)\log\epsilon^{-1}\right)$.